# Conceptualization of this study, Methodology, Software 2]organization=Academy of Mathematics and Systems Science, Chinese Academy of Sciences, city=Beijing, postcode=100190, country=China

### v1

@@ 4.2 Real world data
However, it’s important to consider that we may not always have access to large datasets, such as the M4 dataset, for training. If we only have a small amount of data, directly extracting features for model training may not yield satisfactory results. To address this situation, we propose two solutions. The first approach involves using a previously trained model. If we have trained the OSP model on another large dataset, we can directly use it to make predictions on the current dataset. Although we advocate for the training data to be similar to the prediction dataset, considering the large volume of data used to train the model in the previous dataset, it can learn to a certain extent to generate corresponding predictions for similar situations in the current dataset’s features. The second method takes into account that our model has relatively flexible requirements for training datasets. We can consider using a generated simulation dataset for training. For instance, GRATIS is capable of efficiently generating new time series with controllable features Kang et al. (2019). This method can effectively augment our original dataset. We can train the model on the newly generated data and then apply it to the actual data to enhance the prediction, a strategy that has proven to be effective.
Table 8: The prediction results of the OSP model via trained M4 data.
@@
Table 9: The prediction results of the OSP model via simulated data from GRATIS.
@@
4.3 Discussion of other content
Forecast combination In the empirical section, we consistently used the interval with the smallest average error as the target for training the model, ultimately generating prediction results for four distinct settings. However, when we have no prior knowledge about the future, choosing one out of the four prediction results can be challenging. Therefore, we suggest considering the method of forecast combination. The most straightforward approach is to average all the forecast results and use them as the final forecast result. Alternatively, we can average the improved prediction results of the classification model or regression model separately and then use them as predictions.
Additionally, we propose that an auxiliary meta-model can be trained to pre-assign weights to each prediction result to obtain the final combined prediction. For instance, when training the meta-model, you can input the features of the training set data, with the objective being to minimize the errors of the four prediction results of the training set data. This allows it to assign different prediction weights to new time series data.
In the table, we output the combined forecast results using the average method in the M4 data forecast results. It can be seen that although the combined forecast results are not as good as the previous optimal forecast, they can still be effectively improved compared to using the total time series. This also resolves the difficulty of not knowing how to choose the final prediction.
Table 10: Prediction error of combination prediction and random starting point.
@@ 4.3 Discussion of other content
ETS Total series 18.06 3.44 12.20 1.15 16.83 0.95
Full model 16.19 2.62 11.67 0.97 16.46 0.85
@@
thetaf Total series 17.06 3.40 11.82 1.22 15.68 0.97
Full model 16.82 2.51 11.79 0.99 15.44 0.86
@@
Verification of Model Effectiveness To further validate the efficacy of the model and to demonstrate that the improvement in prediction is not merely due to time series interception, we compared the OSP model with a random strategy. In the random strategy, we randomly select an interval as the starting interval, then select ‘n’ equally divided points as the starting points of the new sequence, construct the new sequence, make predictions, and then average the prediction results. The prediction results of the random strategy are presented in the table. It can be observed that our OSP model significantly outperforms the random strategy. Concurrently, the table also indicates that incorrect interception of data for prediction can lead to a decrease in prediction accuracy.
@@ 5 Conclusions and future work
This paper presents a novel OSP-TSP approach for optimizing the forecast performance by accurately determining the optimal starting point of the input time series. To simplify the model training, we narrow the prediction goal to focus on predicting the optimal forecast starting interval. Specifically, the training data is evenly divided into multiple sub-intervals, where the prediction errors at equally spaced points within each sub-interval are calculated and recorded. The interval containing the optimal starting point is then identified and marked. We employ XGBoost and LightGBM models to predict this optimal starting interval. In the final prediction step, predictions are generated from a few selected, evenly distributed points within the identified interval. These predictions are used to construct tailored sequences for the baseline model. By averaging the outputs, we achieve a final forecast that reflects enhanced accuracy.
To demonstrate the effectiveness of the proposed method, we conducted several empirical analysis using the M4 dataset as well as several other real-world datasets. The results indicate that forecasts utilizing optimal starting points yield greater accuracy compared to those baseline models which use the complete data. Moreover, with respect to improvement, applying the optimal starting interval significantly reduces target prediction errors while effectively mitigating other types of prediction errors.
@@
Furthermore, we propose two approaches for addressing data scarcity issue. The first approach involves making predictions with a pre-trained model, which can fully leverage the knowledge and experience of the existing model. The second approach uses a generated simulation datasets for training. For example, tools like GRATIS can be employed to generate simulated time series with controllable characteristics for effectively augmenting the original datasets. Models trained on these simulated datasets can then be applied to real-world data so that the prediction accuracy can be effectively enhanced.
@@
We conclude this article with several interesting future topics. First, this study lacks an in-depth analysis of the specific selection criteria for multiple improved results provided by the OSP approach. It is of great interest to train an additional model, which can assign weights to the different improved forecasts for a even better prediction performance. Moreover, this paper only conducts empirical research using ETS and thetaf as the baseline models. In theory, the choice of baseline models can be more diverse and advanced. Future work could involve a more comprehensive empirical analysis along this direction, exploring a wider range of baseline models to discover additional opportunities for enhancing prediction accuracy.
Acknowledgements
This research is supported by the National Social Science Fund of China (22BTJ028)

### v2

@@ 4.2 Other real world data
It is worth noting that the practitioners may not always have access to large datasets, such as the M4 dataset, for training. If we only have a small amount of data, directly extracting features for model training might not yield satisfactory results. To address this issue, we propose two approaches. The first approach involves using a pre-trained model. If we have trained the OSP model on another large dataset, we can directly use it to make predictions on the current dataset. While we advocate for the training data to closely resemble the prediction dataset, given the substantial volume of data employed in training the model in our previous experiments, it has demonstrated the ability to learn and generate corresponding predictions for analogous situations present in the features of the current dataset. The second approach is to generate a simulation dataset for training, noticing that our model has relatively flexible requirements for training datasets. For instance, the GRATIS is capable of efficiently generating new time series with controllable features Kang et al. (2019). This method can effectively augment our original dataset. We can train the model on the newly generated data and then apply it to the actual data to enhance the prediction, a strategy that has proven to be effective. We have developed two numerical experiments to demonstrate the effectiveness of the proposed two approaches. The prediction error results are displayed in Table 8 and Table 9, respectively. According to 8 and Table 9, we find that the MASE error results obtained using our pre-trained OSP-TSP method are smaller in almost all cases as compared to those prediction results obtained by using the total time series. This indicates that our method indeed yields better prediction results as well as our pre-trained approaches. The MAPE results did not outperform those obtained using the total time series in some cases, which is reasonable since our method does not use MAPE as the target error to be minimized.
Table 8: Prediction error results of five real world datasets via pre-trained OSP model on M4 data.
@@
Table 9: Prediction error results of five real world datasets via pre-trained OSP model on simulated data using GRATIS.
@@
4.3 Extensive experiments
Forecast combination. In the empirical section, we use the interval with the smallest average error as the target for training the model, ultimately generating prediction results for four distinct settings. However, when we have no prior knowledge about the future, choosing one out of the four prediction results can be challenging. Therefore, we suggest considering the method of forecast combination. The most straightforward approach is to average all the forecast results and use them as the final forecast result. Alternatively, we can average the improved prediction results of the classification model or regression model separately and then use them as predictions. Moreover, we propose that an auxiliary meta-model can be trained to pre-assign weights to each prediction result to obtain the final combined prediction. For instance, when training the meta-model, you can input the features of the training set data, with the objective being to minimize the errors of the four prediction results of the training set data.
In Table 10, we report the combined prediction results using the average method for yearly, quarterly and monthly M4 data. It can be seen that although the combined forecast results are not as good as the previous optimal prediction results, they can still be effectively improved as compared to those using the total time series. This also resolves the difficulty of not knowing how to choose the final prediction. This also provides a viable approach for selecting the final prediction results.
Table 10: Prediction error results of combined prediction and random starting point.
@@ 4.3 Extensive experiments
ETS Full model 16.19 2.62 11.67 0.97 16.46 0.85
@@
Total series 18.06 3.44 12.20 1.15 16.83 0.95
thetaf Full model 16.82 2.51 11.79 0.99 15.44 0.86
@@
Total series 17.06 3.40 11.82 1.22 15.68 0.97
Extensive results on machine learning models. Note that the ETS and thetaf models already give more weight to the last observations, the proposed method might not offer significant improvements when the base model performs well on the total time series. Therefore, we further consider using the machine learning models to illustrate the advantages of the proposed method, where in principle all instances are weighted equally without weighting scheme. Specifically, we consider neural network autoregressive method (nnetar) provided in the forecast package as an equally weighted approach for prediction. The experiment is conducted on M4 data. The prediction error results are displayed in Table 11. According to Table 11, We find that the prediction MASE results of the proposed OSP-TSP method outperform those derived from total time series under the same machine learning base model. And the MAPE results lead to similar conclusions. These findings highlight the effectiveness of our approach in refining prediction accuracy. Although we observe that the MAPE results on hourly data are not better than those obtained from using the total time series. It is reasonable since our optimization target error is MASE rather than MAPE.
Table 11: Prediction performance of neural network autoregressive method.
Category Yearly Quarterly Monthly
MAPE MASE MAPE MASE MAPE MASE
Actual mean 20.82 3.54 16.50 1.35 19.01 1.03
Average mean 21.21 3.57 15.38 1.37 18.64 1.03
Total series 21.61 4.06 15.89 1.57 19.80 1.15
Category Weekly Daily Hourly
MAPE MASE MAPE MASE MAPE MASE
Actual mean 8.26 2.40 4.82 3.54 25.98 1.03
Average mean 8.50 2.56 5.08 3.42 23.82 1.01
Total series 9.91 3.86 5.82 4.08 14.67 1.08
Comparison with change point detection methods. We consider using the cpm package in R (Documentation, 2024) to apply change point detection technique in our comparison experiment. It contains several change-point detection methods and we adopt the CUSUM test provided by the cpm package, which is one of the most commonly used change-point detection method (Healy, 1987; Bücher et al., 2019). Specifically, we first adopt the change-point detection method based on CUSUM provided to identify potential structural breaks within the data. Once the change-point is detected, we then extract the subsequence starting from the identified change-point to the end of the series. This subsequence is then used for model training and forecasting analysis. The experiment is conducted on M4 data. The prediction error results are displayed in Table 12. We find that the prediction performances of the proposed OSP-TSP method outperform those derived from subsequences based on change-point detection. This may be because the purpose of change point detection methods is to seek structural changes within a sequence, rather than to achieve the best prediction results. Therefore, relying solely on change point detection methods to find the optimal subsequence sequence for forecasting might not be satisfactory.
Table 12: Prediction error results of OSP-TSP method against change point detection method on M4 data.
Model Yearly Quarterly Monthly
MAPE MASE MAPE MASE MAPE MASE
OSP best 16.39 2.42 11.75 0.94 15.49 0.83
cpm-Thetaf 16.90 2.97 12.04 1.18 15.91 0.97
cpm-ETS 17.68 2.98 12.03 1.12 17.11 0.95
Model Weekly Daily Hourly
MAPE MASE MAPE MASE MAPE MASE
OSP best 9.11 1.98 4.81 2.69 21.65 1.61
cpm-Thetaf 9.19 2.67 5.62 3.31 23.81 2.48
cpm-ETS 9.34 9.34 5.28 3.31 29.44 1.75
Combined with the FFORMA model. As we mentioned previously, the proposed OSP-TSP approach is a general framework, which can be integrated with a variety of time series forecasting models. Then it is of great interest to study whether our OSP-TSP approach can enhance the performance of some state-of-the-arts forecasting models. For example, the FFORMA (Montero-Manso et al., 2021) model, which achieved the second place in the M4 competition. We consider using the FFORMA model as our base model to compare the results with and without the application of the our OSP-TSP approach. Specifically, we use the M4 weekly data to illustrate the finite sample performance. For the experimental setup, 30% of the data is used to train the FFORMA model, while 40% is employed to train the optimal prediction starting point model. The remaining 30% of the data is used to evaluate the improvement results of the proposed method. The outputs include prediction improvements obtained by training the model with both the actual optimal starting point and the averaged optimal starting point as targets. The prediction error results are reported in Table 13. These results encompass predictions from sub-models representing specific starting point strategies, as well as the aggregated results for the two major categories (actual and averaged optimal starting points). The evaluation results show that nearly all predictions employing optimal starting points achieve improvements in MASE. However, the performance in MAPE is relatively less consistent, likely due to the fact that the training objective was based on MASE. This suggests that while the proposed method effectively enhances the accuracy of predictions in terms of scale-invariant measures, further refinements may be needed to optimize performance for other error metrics such as MAPE.
Table 13: Prediction error results on weekly data of M4 with base model FFORMA.
Class Method MAPE MASE
Actual xgbcls 9.82 1.66
xgbreg 10.30 1.64
lgbreg 10.07 2.00
lgbcls 10.48 1.72
Average xgbcls 9.65 1.62
xgbreg 10.08 1.61
lgbreg 9.95 1.95
lgbcls 10.01 1.67
Actual mean 10.00 1.82
Average mean 9.76 1.81
Total series 9.42 2.00
@@ 5 Conclusions and future work
In this paper, we propose a novel OSP-TSP approach for optimal forecasting. Specifically, the proposed approach contains two main steps. In the first step, the training data is evenly divided into multiple sub-intervals. We then employ XGBoost and LightGBM models to predict the optimal starting interval, and the interval that contains the OSP is then identified. In the second step, predictions are generated from a few selected, evenly distributed points within the identified interval. These predictions are used to construct tailored sequences for the baseline model. By averaging the outputs, we achieve a final prediction result that reflects enhanced accuracy.
To demonstrate the effectiveness of the proposed method, we conducted several empirical analysis on the M4 dataset as well as several other real world datasets. The results indicate that forecasts utilizing optimal starting points yield better prediction performances as compared to those baseline models which use the total time series. We find that the proposed OPS-TSP approach can be particularly useful, when the time series data have potential structural breaks or concept drifts. This is because when time series suffer from potential structural changes or concept drift, using the complete sequence for forecasting typically fails to achieve good results. Our proposed method, however, can automatically select the subsequence that achieves the optimal forecasting performance, which also can overcome the challenges posed by structural changes or concept drifts. For instance, the stock data during the implementation of specific financial policies (Andreou and Ghysels, 2009; Mahata et al., 2020), and the tourism industry data before and after the COVID-19 pandemic (Kourentzes et al., 2021; Wu et al., 2023).
@@
To address the data scarcity issue, we further propose approaches that making predictions with a pre-trained OSP model, which can fully leverage the knowledge and experience of the existing method. For example, we can use an OSP model pre-trained on M4 data. Moreover, tools like GRATIS can be employed to generate simulated time series with controllable characteristics for effectively augmenting the original datasets. Models trained on these simulated datasets can then be applied to real-world data so that the prediction accuracy can be effectively enhanced.
@@
We conclude this article with several interesting future topics. First, the proposed OSP-TSP approach involves multiple forecasting and sub-segmentation, which might be computationally demanding for larger datasets or high-frequency data. Then how to reduce the computational costs of the OSP-TSP approach becomes a problem of interest. For instance, since the calculation for subsequence are independent, a parallel computing scheme can be directly applied to reduce the computation time. Moreover, we can apply some heuristic sorting algorithms or adopt the feature based forecasting algorithms (Talagala et al., 2022) to reduce the computational costs. Second, this study lacks an in-depth analysis of the specific selection criteria for multiple improved results provided by the OSP approach. It is of great interest to train an additional model, which can assign weights to the different improved forecasts for a even better prediction performance. Third, this paper only conducts empirical research using a few forecasting models as the baseline models. In theory, the choice of baseline models can be more diverse and advanced. Future work should involve a more comprehensive empirical analysis along this direction, exploring a wider range of baseline models to discover additional opportunities for enhancing prediction accuracy.